# Can Automated Feedback Turn Students into Happy Prologians?

### v1

@@
Abstract
Giving personalized feedback to students is very important to the learning process. However, doing so in a timely manner can be difficult to accomplish in very large courses. Recent work has explored different types of automated feedback adapted to different languages and programming paradigms, particularly logic programming. In ProHelp, we implemented several of these types of feedback so that they could be used by students enrolled in a logic programming class. Then, we surveyed those students to find if the feedback was useful and which types of feedback they preferred. Results show that students found all types of feedback helpful, with automatic testing, in particular, being the most helpful type. We also explore student preferences for which types of feedback they would most like to see implemented in the future.
Keywords:
Automated Feedback Logic Programming User Survey Computer-aided Education.
1 Introduction
Giving timely and formative feedback to students is very important for the learning process [11, 28]. However, doing so in large courses is a difficult task [19], and even impossible in cases like Massive Open Online Courses (MOOCs). Over the years, many automated assessment frameworks have been proposed to address this problem. Automated assessment tools such as Mooshak/Enki [20] or GitSEED [18], among others [6, 10, 22, 26], allow faculty to automatically test students’ submissions and provide them with varying levels of feedback.
@@ 1 Introduction
Novice students usually find it hard to debug logic programs. This is mainly due to the lack of traditional control and data flows they learn in imperative languages. Hence, we propose ProHelp, an automated feedback tool for Prolog that can be integrated with GitSEED. ProHelp includes feedback features that range from automatic testing to typo correction suggestions and validation of implementation techniques.
@@
The two main goals of this work are to (1) understand which types of feedback implemented in ProHelp students find the most helpful, and (2) gauge students’ interest in more complex types of automated feedback they would like to see in the future. To this end, we designed and deployed a survey with the students that interacted with ProHelp. Results show that the feedback provided to students was useful and the system has good usability. Students also showed great interest in proposed improvements and suggested some themselves.
2 Related Work
2.0.1 Automated Assessment Tools (AATs).
@@ 2 Related Work
Automated assessment of programming assignments has a long history of more than 60 years of research [19]. Most Automated Assessment Tools assess programs by either comparing them with a known-correct solution or by executing a series of tests and checking their results [11]. A large number of current AATs are based around Integrated Development Environment-like web-interfaces with examples like CodeOcean [26], Mooshak/Enki [14, 20] and Web-Cat [6]. Other platforms, like ProgEdu [3], GitHub Classroom [^1] https://classroom.github.com/ and GitSEED [18] focus on Git, taking advantage of Continuous Integration (CI) actions to assess students’ work as it is submitted to the repository. These tools can be easier for students to use because they are often already familiar with the Git workflow.
2.0.2 Automated Feedback for Logic Programming.
@@
Many systems have been proposed for automated feedback, debugging and repair of logic programming languages [13, 16, 27, 1]. These tools face different challenges compared with other programming paradigms due to the declarative nature of the languages, lack of explicit control flow and issues like non-determinism, non-monotonic reasoning or non-termination [16]. INCOM [13] introduces a two-stage tutoring model where students first engage in task analysis progressing to the implementation of Prolog predicates. This structured approach assists learners in bridging the gap between problem specification and algorithm development. Meanwhile, PRAM [16] focuses on the automatic assessment of Prolog programs by incorporating both static and dynamic metrics to evaluate code style, complexity, and dynamic correctness, thereby reducing instructor workload and facilitating the shift from procedural to declarative thinking. Complementing these educational tools, ProFL [27] applies fault localization techniques, particularly coverage-based and mutation-based approaches, to identify errors in Prolog programs effectively. Moreover, the FormHe [1] system extends automated feedback to Answer Set Programming by combining traditional fault localization methods with modern strategies like Large Language Model (LLM)-based fault localization and repair, and mutation-based program repair techniques.
@@
One common element of earlier works on automated logic programming feedback is the use of structured, rule based feedback, which can force students to solve problems in a certain way. Meanwhile, more recent approaches follow a testing-based fault localization method which helps students find problems in their programs instead of guiding them in a particular direction. However, these newer tools lack comprehensive user studies to evaluate their performance. Our work addresses this research gap by directly engaging with the target audience of the tool – the students – and prioritizing their needs.
3 Automated Assessment System
At the core of this work is ProHelp, a new tool for automatic feedback to students learning logic programming with Prolog. ProHelp was integrated with GitSEED [18], a recent easy-to-use automated assessment tool based on GitLab. In this section, we present the overall framework, as well as the feedback extensions we implemented on top of GitSEED.
3.1 GitSEED
Figure 1: Simplified overview of GitSEED’s operation scheme.
GitSEED [18] is an open-source, language-agnostic automated assessment tool integrated with GitLab. It provides students with personalized feedback on programming assignments while they learn Git fundamentals. Unlike traditional assessment platforms, GitSEED operates within GitLab’s CI workflow, eliminating the need for students to learn new interfaces.
@@ 3.1 GitSEED
As Figure 1 shows, when students submit their work to GitLab, GitSEED automatically evaluates it using a predefined test suite and program analysis tools specified by the faculty. The evaluation report is then pushed directly to the students’ git repositories, ensuring immediate access to personalized feedback.
3.2 Automatic feedback for Prolog submissions
@@
Figure 2: Screenshots of the different types of student feedback available in ProHelp.
The minimum configuration of GitSEED simply runs a test suite on the student’s submission and reports the results as feedback. However, as shown in Figure 1, GitSEED can be extended with custom program analysis tools that give students more personalized feedback. Taking advantage of this customizability, we implemented several types of feedback in ProHelp that are specific to the Prolog context. Below, we describe each of the 7 types of automated feedback, along with an example for each one, shown in Figure 2.
@@
3.3 Future directions
One of the main goals of this work is to gauge students’ opinions on the feedback features already implemented to decide which ones to prioritize for future work. Planned features include several AI-based procedures using formal methods targeted to the specific context of logic programming.
@@ 3.3 Future directions
  • Test output differences:  The test suites in ProHelp are composed of unit tests implemented through plunit. Almost always, these tests consist of a series of predicate calls followed by an assertion that compares a value with the expected response. It would be useful for students to provide more information when these assertions fail, particularly showing the difference between generated and expected values.
4 Methodology
ProHelp was deployed in a BSc-level class on logic reasoning and programming with 365 students at University X [^2] Anonymized for double-blind review.. During the course, students were given 3 different types of Prolog assignments: a 5-part story-driven introductory exercise, recitation exercises and a course project. Of these, the first two types were optional and ungraded, while the project was mandatory and represented 50% of the final grade of the course.
Overall, 312 students interacted with ProHelp during the course. We obtained 148 responses to our survey (47%), of which we dropped 4 due to failed attention checks [17]. Of the 144 valid responses, 28 respondents identified as female (19%), 111 as male (77%), 1 as other (<1%) and 4 preferred not to disclose (3%). While we have no information about the gender distribution for students who took this class, the official distribution for the enrolled students in this BSc degree is 83% male students and 17% female. This provides some confidence that our sample is representative of the overall course population.
To understand the strengths and shortcomings of ProHelp, we performed a post-experience survey. The survey had an estimated duration of 15 minutes and was performed using EUSurvey [^3] https://ec.europa.eu/eusurvey, during the two weeks following the last programming assignment. With this survey, we plan to answer the following three research questions:
@@ 4.1 Data Analysis
Our study focuses on analyzing 3- and 5-point Likert scales regarding the usefulness of different aspects of the feedback received. For RQ1 and RQ3, we are interested in how different answers of the same student compare to one another (the ratings for the different types of feedback). To accomplish this, we use the Friedman rank sum test [7], to check for any statistically significant differences between the feedback types. In cases where there are such differences, we employ the post-hoc pairwise Durbin-Conover test [4] with the Holm correction method [9], in order to identify which pairs of feedback types have statistically significant differences in usefulness.
@@
In RQ2, we are interested in how different sub-groups of students rated the usefulness of ProHelp (on a 5-point Likert scale). To accomplish this, we use different tests depending on the nature of the independent variable. Students’ interest was measured on a scale from 1 to 5 (an ordinal variable) and, as such, we use the Spearman’s rank correlation [25] to compare it with the usefulness. Meanwhile, students’ engagement and usage of LLMs were both measured through categorical variables. We use the Kruskal-Wallis test [8] for the engagement since it has more than 2 categories, and the Mann-Whitney $U$ test [15] for the LLM-usage since it only has 2. The significance level for all tests was set at $0.05$.
@@
Finally, our survey also contained two open-ended questions regarding future feedback suggestions and general opinions on the system. The answers were coded according to a unified code-book for Feedback and Improvements. The code-book was created by author X. The answers were coded, independently, by authors X and Y, with any differences conciliated through a posterior discussion.
4.2 Ethics and Privacy
@@
All students had access to the same system with the same types of feedback available. This ensured that no group of students had any advantage over the rest. Participation in this study was voluntary, unremunerated, and all participants provided informed consent for data collection and processing before the study. Participants were recruited through a course announcement after the end of the last assignment. No identifiable personal data was collected. In University X, where the course was located and the survey performed, studies with these characteristics do not need to be submitted to the Ethics Committee (as confirmed by the Chair of the Ethics Committee). Co-authors from other institutions than University X only had access to aggregate statistics and not individual responses. We have completed the European Commission’s ethics self-assessment checklists and comply with all the guidelines.
5 Results
@@
This work aims to understand the advantages and limitations of the ProHelp system and explore what types of feedback students would like to have in the future. The statistical analysis and plots presented in this section were created using R version 4.4.2 [23], statsExpressions [21], ggstats [12] and ggsignif [5].
@@
Figure 3: Agreement distribution for the statement “I think the automated feedback helped increase my grade in this course.”.
5.1 Usefulness and usability of the feedback (RQ1)
@@
To gauge if ProHelp was useful to students, we asked them if they agreed with the statement “I think the automated feedback helped increase my grade in this course” using a 5-point Likert scale. Figure 3 shows that 74% of student responses agreed that the provided feedback was useful, while 11% disagreed and 15% neither agreed nor disagreed. This indicates that ProHelp was helpful for students. We also performed a usability analysis through the System Usability Scale (SUS) questionnaire [^4] We used a peer reviewed translation of the SUS questions. The language and reference have been omitted for anonymity and will be included in the final version. [2]. The average SUS score among our respondents was 78.5 which corresponds to a B+ grade according to the curved grading interpretation by Sauro and Lewis [24]. While there is room for improvement, the results show that our system is fairly usable in its current state.
@@
Figure 4: Usefulness rating of the different types of feedback. The median response for each type is shown by the vertical line. The right side of the figure shows pairs of feedback types for which the Durbin-Conover post-hoc test indicates statistically significant differences in ratings. For example, there is a statistically significant difference between the responses for “Automatic testing” and “Score rankings” with $p=1.23\mathrm{e}\text{-}07$.
@@
This study also aims to understand if any type of feedback stood out as particularly more useful than the others. Hence, we asked the students to rank each of the 7 types of feedback from Section 3.2 from 1 (not useful) to 5 (very useful). Figure 4 presents the answer distribution to these questions. Most students ranked all 7 types of feedback as useful, with all of the medians being $\geq 4$.
@@
Most students consider that the provided feedback was helpful and the system has a usability score of B+. The most useful feedback were “Automatic testing”, “Open choice points warning” and “Predicate scoring”, while the least useful were “Score ranking” and “Solution type validation”. These conclusions are supported by statistically significant differences.
5.2 Impact of interest, engagement and LLM use (RQ2)

### v2

@@
Abstract.
Providing valuable and personalized feedback is essential for effective learning, but delivering it promptly can be challenging in large-scale courses. Recent research has explored automated feedback mechanisms across various programming languages and paradigms, including logic programming.
In this work, we present a student survey were we evaluate the perceived usefulness of different feedback types and identified which are most valued. Our results indicate that students found all implemented feedback types helpful, with automatic testing ranked as the most useful. We also introduce a dataset comprising 7201 correct and incorrect Prolog submissions, along with 200 manually annotated programs labeled with bug types and corresponding corrections. Finally, we explore student preferences for which types of feedback they would most like to see implemented in the future.
Automated Feedback, Logic Programming, User Survey, Dataset, Computer-aided Education
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Applied computing Computer-assisted instructionccs: Human-centered computing Empirical studies in HCIccs: Theory of computation Constraint and logic programming
1. Introduction
Giving timely and formative feedback to students is very important for the learning process (Keuning et al., 2019; Wisniewski et al., 2020). However, doing so in large courses is a difficult task (Paiva et al., 2022), and even impossible in cases like Massive Open Online Courses (MOOCs). Over the years, many automated assessment frameworks have been proposed to address this problem. Automated assessment tools such as Mooshak/Enki (Paiva et al., 2016) or GitSEED (Orvalho et al., 2024), among others (Edwards and Pérez-Quiñones, 2008; Iddon et al., 2023; Peveler et al., 2017; Staubitz et al., 2016), allow faculty to automatically test students’ submissions and provide them with varying levels of feedback.
@@ 1. Introduction
However, despite these advances, significant challenges remain in the context of logic programming. Novice students usually find it hard to debug logic programs. This is mainly due to the lack of traditional control and data flows they learn in imperative languages. In this work, we aim to better understand these difficulties and students’ behavior when learning Prolog. To accomplish this, we created ProHelp, an automated assessment platform for Prolog built on top of GitSEED. ProHelp includes feedback features that range from automatic testing to typo correction suggestions and validation of implementation techniques.
We designed and deployed a mixed-methods study with the students that interacted with ProHelp. We used survey responses and submission logs collected during a 9-week undergraduate course with 365 students. Overall, we analyzed 7201 submissions and 144 survey responses.
The main contributions of this work are:
  • Understanding which types of feedback implemented in ProHelp students find the most helpful;
  • Understanding what types of bugs students have in their submissions;
  • A dataset of 7201 correct and incorrect Prolog programs, with 200 programs manually annotated with bug types and corrections;
  • Gauging students’ interest in more complex types of automated feedback they would like to see in the future.
@@
Results show that the feedback provided to students was helpful, and that the automated assessment system has good usability. Students also showed great interest in proposed improvements and suggested some themselves.
This paper is organized as follows: in Section 2 we present an overview of related work. Following that, in Section 3, we present the course during which this study was performed and in Section 4 describe the automated assessment system and the types of feedback provided. Then, in Section 5, we present the methodology for our study and in Section 6, we present the analysis and results. Finally, in Section 7, we present a discussion on the obtained results and conclude the paper.
2. Related Work
Automated Assessment Tools (AATs).
@@ 2. Related Work
Automated assessment of programming assignments has a long history of more than 60 years of research (Paiva et al., 2022). Most Automated Assessment Tools assess programs by either comparing them with a known-correct solution or by executing a series of tests and checking their results (Keuning et al., 2019). A large number of current AATs are based around Integrated Development Environment-like web-interfaces with examples like CodeOcean (Staubitz et al., 2016), Mooshak/Enki (Leal and Silva, 2003; Paiva et al., 2016) and Web-Cat (Edwards and Pérez-Quiñones, 2008). Other platforms, like ProgEdu (Chen et al., 2017), GitHub Classroom [^1] https://classroom.github.com/ and GitSEED (Orvalho et al., 2024) focus on Git, taking advantage of Continuous Integration (CI) actions to assess students’ work as it is submitted to the repository. These tools can be easier for students to use because they are often already familiar with the Git workflow.
Automated Feedback for Logic Programming.
@@
Many systems have been proposed for automated feedback, debugging and repair of logic programming languages (Le and Pinkwart, 2011; Mansouri et al., 1998; Thompson and Sullivan, 2020; Brancas et al., 2025a). These tools face different challenges compared with other programming paradigms due to the declarative nature of the languages, lack of explicit control flow and issues like non-determinism, non-monotonic reasoning or non-termination (Mansouri et al., 1998). INCOM (Le and Pinkwart, 2011) introduces a two-stage tutoring model where students first engage in task analysis progressing to the implementation of Prolog predicates. This structured approach assists learners in bridging the gap between problem specification and algorithm development. Meanwhile, PRAM (Mansouri et al., 1998) focuses on the automatic assessment of Prolog programs by incorporating both static and dynamic metrics to evaluate code style, complexity, and dynamic correctness, thereby reducing instructor workload and facilitating the shift from procedural to declarative thinking. Complementing these educational tools, ProFL (Thompson and Sullivan, 2020) applies fault localization techniques, particularly coverage-based and mutation-based approaches, to identify errors in Prolog programs effectively. Moreover, the FormHe (Brancas et al., 2025a) system extends automated feedback to Answer Set Programming by combining traditional fault localization methods with modern strategies like Large Language Model (LLM)-based fault localization and repair, and mutation-based program repair techniques.
@@
One common element of earlier works on automated logic programming feedback is the use of structured, rule-based feedback, which can force students to solve problems in a certain way. Meanwhile, more recent approaches follow a testing-based fault localization method which helps students find problems in their programs instead of guiding them in a particular direction. However, these newer tools lack comprehensive user studies to evaluate their performance. Our work addresses this research gap by directly engaging with the target audience of the tool – the students – and prioritizing their needs.
3. Course Structure
We evaluated student’s behavior during a 9 week-long bachelor’s level logic programming course. During this course, students learned classic logic reasoning, as well as logic programming through Prolog. Overall, students faced three different types of Prolog exercises during this course: a 5-part role-playing exercise to familiarize students with the language, 22 optional recitation exercises, and one graded and mandatory final project. The mandatory project was worth 50% of the final grade, while the remaining 50% was from other theoretical assessments. Next, we describe the Prolog exercises in more detail.
  • Role-playing Exercise
    • Optional, not graded, online only
    • 5 puzzles over 5 days
    • A series of simple logic puzzles where students help a detective catch a group of criminals using logic programming;
  • Recitation Exercises
    • Optional, not graded, partially solved in-class using pen and paper
    • 22 exercises
    • 4 Prolog exercise sheets grouped into the topics of: lists, arithmetic, negation and higher-order functions;
  • Project
    • Mandatory, graded
    • 5 weeks duration
    • Worth 50% of the final course grade. The goal of this year’s project was to create a solver for the logic puzzle star battle;
[…]
Figure 1. Course exercises timeline.
In Figure 1, we show the timeline of the different exercises solved by the students. During the first weeks, the students learned logic fundamentals. Then, during week 3, the students started learning Prolog, with the optional exercise sets being released during weeks 3, 4 and 5. Meanwhile, the project statement was released at the beginning of week 4 and lasted until the end of week 9.
4. ProHelp: Prolog Automated Assessment Platform
Figure 2. Simplified overview of GitSEED’s operation scheme.
@@ 4. ProHelp: Prolog Automated Assessment Platform
The submission and evaluation platform used in this course was ProHelp, an extension built on top of the open-source automated assessment framework GitSEED (Orvalho et al., 2024). GitSEED provides students with personalized feedback on programming assignments while they learn Git fundamentals. Unlike traditional assessment platforms, GitSEED operates within GitLab’s CI workflow, eliminating the need for students to learn new interfaces. As Figure 2 shows, when students submit their work to GitLab, GitSEED automatically evaluates it using a predefined test suite and program analysis tools specified by the faculty. The evaluation report is then pushed directly to the students’ git repositories, ensuring immediate access to personalized feedback.
@@
Figure 3. Screenshots of the different types of student feedback available in ProHelp.
4.1. Automated Feedback
The minimum configuration of GitSEED simply runs a test suite on the student’s submission and reports the results as feedback. However, as shown in Figure 2, GitSEED can be extended with custom program analysis tools that give students more personalized feedback. Taking advantage of this customizability, we created ProHelp and implemented seven types of feedback that are specific to the Prolog context. Below, we describe each of the 7 types of automated feedback, along with an example for each one, shown in Figure 3.
@@ 4.1. Automated Feedback
4.2. Future Directions
One of the goals of this work is to gauge students’ opinions on the feedback features already implemented and decide which ones to prioritize for future work. Planned features include several AI-based procedures using formal methods targeted to the specific context of logic programming.
@@ 4.2. Future Directions
  • Test output differences:  The test suites in ProHelp are composed of unit tests implemented through plunit. Almost always, these tests consist of a series of predicate calls followed by an assertion that compares a value with the expected response. It could be useful for students to provide more information when these assertions fail, particularly showing the difference between generated and expected values.
5. Methodology
ProHelp was deployed in a BSc-level class on logic reasoning and programming with 365 students at Universidade de Lisboa. During the course, students were given automated feedback on 3 different types of Prolog assignments: a role-playing exercise, recitation exercises and a final project.
To understand the strengths and shortcomings of ProHelp, we performed a post-experience survey. The survey, which is replicated in Appendix A, had an estimated duration of 15 minutes and was performed using EUSurvey [^2] https://ec.europa.eu/eusurvey during the two weeks following the last programming assignment.
Of the 312 students that interacted with ProHelp during the course, we obtained 148 responses to our survey (47%), of which we dropped 4 due to failed attention checks (Oppenheimer et al., 2009). Of the 144 valid responses, 28 respondents identified as female (19%), 111 as male (77%), 1 as other (¡1%) and 4 preferred not to disclose (3%). While we have no information about the gender distribution for students who took this class, the official distribution for the enrolled students in this BSc degree is 83% male students and 17% female. This provides some confidence that our sample is representative of the overall course population.
In this work, we plan to answer the following four research questions:
@@ 5. Methodology
    What types of bugs did students introduce in their Prolog programs?
  4. RQ4.
@@ 5.1. Data Analysis
Part of our study focuses on analyzing 3- and 5-point Likert scales regarding the usefulness of different aspects of the feedback received. For RQ1 and RQ4, we are interested in how different answers of the same student compare to one another (the ratings for the different types of feedback). To accomplish this, we use the Friedman rank sum test (Friedman, 1937), to check for any statistically significant differences between the feedback types. In cases where there are such differences, we employ the post-hoc pairwise Durbin-Conover test (Conover, 1999) with the Holm correction method (Holm, 1979), in order to identify which pairs of feedback types have statistically significant differences in usefulness.
@@
In RQ2, we are interested in how different sub-groups of students rated the usefulness of ProHelp (on a 5-point Likert scale). To accomplish this, we use different tests depending on the nature of the independent variable. Students’ interest was measured on a scale from 1 to 5 (an ordinal variable) and, as such, we use the Spearman’s rank correlation (Spearman, 1904) to compare it with the usefulness. Meanwhile, students’ engagement and usage of LLMs were both measured through categorical variables. We use the Kruskal-Wallis test (Hollander et al., 2013) for the engagement since it has more than 2 categories, and the Mann-Whitney $U$ test (Mann and Whitney, 1947) for the LLM-usage since it only has 2. The significance level for all tests was set at $0.05$.
@@
The survey also contained two open-ended questions regarding future feedback suggestions and general opinions on the system. The answers were coded according to a unified code-book for Feedback and Improvements. The code-book was created by the first author. The answers were coded, independently, by the first and second authors, with any differences conciliated through a posterior discussion.
At the end of the course, students were given the option to consent to the use of their code for research purposes. The submissions from students who provided consent were fully anonymized prior to any analysis or code inspection. A subset of these submissions was manually examined to categorize the types of bugs students introduced. The buggy programs were labeled using a hierarchical, emergent coding scheme created by the first author.
5.2. Ethics and Privacy
@@
All students had access to the same system with the same types of feedback available. This ensured that no group of students had any advantage over the rest. Participation in the survey was voluntary, unremunerated, and all participants provided informed consent for data collection and processing before the survey. Participants were recruited through a course announcement after the end of the last assignment. No identifiable personal data was collected. In Universidade de Lisboa, where the course was located and the survey performed, studies with these characteristics do not need to be submitted to the Ethics Committee (as confirmed by the Chair of the Ethics Committee). Co-authors from other institutions than Universidade de Lisboa only had access to aggregate statistics and not individual responses. We have completed the European Commission’s ethics self-assessment checklists and comply with all the guidelines.
6. Results
@@
This work aims to understand the advantages and limitations of the ProHelp system and explore what types of feedback students would like to have in the future. The statistical analysis and plots presented in this section were created using R version 4.4.2 (R Core Team, 2024), statsExpressions (Patil, 2021), ggstats (Larmarange, 2025) and ggsignif (Constantin and Patil, 2021).
@@
Figure 4. Agreement distribution for the statement “I think the automated feedback helped increase my grade in this course.”.
6.1. Usefulness and usability of the feedback (RQ1)
@@
To gauge if ProHelp was useful to students, we asked them if they agreed with the statement “I think the automated feedback helped increase my grade in this course” using a 5-point Likert scale. Figure 4 shows that 74% of student responses agreed that the provided feedback was useful, while 11% disagreed and 15% neither agreed nor disagreed. This indicates that ProHelp was helpful for students. We also performed a usability analysis through the System Usability Scale (SUS) questionnaire [^3] We used a peer reviewed translation of the SUS questions. The language and reference have been omitted for anonymity and will be included in the final version. (Brooke, 1996). The average SUS score among our respondents was 78.5 which corresponds to a B+ grade according to the curved grading interpretation by Sauro and Lewis (Sauro and Lewis, 2016). While there is room for improvement, the results show that our system is fairly usable in its current state.
@@
Figure 5. Usefulness rating of the different types of feedback. The median response for each type is shown by the vertical line. The right side of the figure shows pairs of feedback types for which the Durbin-Conover post-hoc test indicates statistically significant differences in ratings. For example, there is a statistically significant difference between the responses for “Automatic testing” and “Score rankings” with $p=1.23\mathrm{e}\text{-}07$.
@@
This study also aims to understand if any type of feedback stood out as particularly more useful than the others. Hence, we asked the students to rank each of the 7 types of feedback from Section 4 from 1 (not useful) to 5 (very useful). Figure 5 presents the answer distribution to these questions. Most students ranked all 7 types of feedback as useful, with all of the medians being $\geq 4$.
@@
A1. Most students consider that the provided feedback was helpful and the system has a usability score of B+. The most useful feedback were “Automatic testing”, “Open choice points warning” and “Predicate scoring”, while the least useful were “Score ranking” and “Solution type validation”. These conclusions are supported by statistically significant differences.
6.2. Impact of interest, engagement and LLM use (RQ2)